# Pair Production in time-dependent Electric field at Finite times

### v5

@@ IV.2 Longitudinal Momentum Spectrum
From Figure 5, we see how particle creation proceeds. At early times, $t=-10[m^{-1}]$ in the QEPP region, when the electric field is increasing, the created particle shows a smooth unimodal-like structure of momentum spectrum with a peak at $p_{\parallel}\sim 2[m].$ The electric field propels the newly created particles towards the negative $z$-direction. The movement of the distribution function’s peak from $p_{\parallel}=2$ $(\leavevmode\nobreak\ eE_{0}\tau)$ to $-2$ $(\leavevmode\nobreak\ -eE_{0}\tau)$ can be understood through the concept of longitudinal quasi-momentum, $P(p_{\parallel},t)$ as depicted in the time evolution shown in Figs. 5(a) to (c). This momentum shift implies that the longitudinal momentum distribution of the generated particles is expected to span a range of $\Delta p_{\parallel}=2$, with the spread determined by the electric field parameters. At $t=0$ (electric field is maximum), the momentum distribution function, $f(p_{\parallel}=0,t=0)\approx 5\times 10^{-3}$, follows the tendency $E^{2}(t)/8$ Schmidt et al. (1999). For $t>0$, a reduction in the field strength with a decrease in an intriguing longitudinal momentum distribution function. It rapidly drops to $99.2\%$ of its maximum value, driven by its reliance on the strength of the electric field at $t=10[m^{-1}]$, as illustrated in Fig. 5(c). Within a narrow range of $-1<p_{\parallel}<1$, the typically smooth unimodal profile of the longitudinal momentum spectra is disrupted, as depicted in Fig. 5(d).
@@
Certainly, we can identify specific time scales associated with the quantum signature that manifests as oscillations within the LMS of created pairs. Drawing from the emergence of a secondary peak, we define three distinct time scales: $(i)$ $t_{cp}$ (central peak formation): This time scale is characterized by the emergence of the secondary peak and the beginning of its development, $(ii)$ $t_{sep}$ (peak separation): At this time, the central peak becomes dominant or the time after the two peaks become distinctly separated, $(iii)$ $t_{dis}$ (disappearance of oscillation): This indicates the time when the oscillations within the central peak fade away or the time after the primary peak (or the left-side peak) ceases to exist. By identifying and quantifying these time scales, we can better characterize the intricate quantum dynamics reflected in the LMS during the REPP stage.
          $E_{0}[E_{c}]$           $t_{cp}[m^{-1}]$           $t_{sep}[m^{-1}]$           $t_{dis.}[m^{-1}]$
@@
Table 2: The effect of electric field strength $E_{0}$ on time scale related to the formation of a central peak.The time scale $t_{cp}$ appearance of a central peak, $t_{sep}[m^{-1}]$ when two peaks become distinctly separated, and $t_{dis.}$ disappearance of the oscillation in LMS.
@@
We have not yet discussed the Keldysh parameter, $\gamma$, which classifies the pair production process as either a multi-photon or tunneling mechanism. Additionally, we are investigating whether the oscillatory behavior explicitly depends on $\gamma$. For this study, we select two different combinations of $E_{0}$ and $\tau$ to vary $\gamma$ and examine its behavior over a finite time interval. Specifically, we set $\gamma=1$, corresponding to two distinct parameter configurations $(E_{0},\tau)$ associated with the Sauter pulse electric field. It is known that $\gamma=1$ represents an intermediate regime involving both tunneling and multi-photon processes, which are inherently non-perturbative Avetissian et al. (2002). This approach allows us to demonstrate how the same $\gamma$ value can lead to qualitatively different outcomes in the time evolution of the pair production process. Figure 6 shows LMS for the Sauter pulse electric field with $E_{0}=0.2E_{c}$ and $\tau=5\,[m^{-1}]$, while Fig.7 shows LMS for field parameter $E_{0}=0.1E_{c}$ and $\tau=10\,[m^{-1}]$. From Figs. 6 (a) to (c) we can see that due to the electric field smooth unimodal peaked profile accelerated towards negative $z-$ direction and after the pulse duration $t=6[m^{-1}]$ we observed smooth profile becomes deformed near the region $-0.5<p_{\parallel}<0.5.$ At $t=11.5[m^{-1}]$, the interference effect is observed, due to which the smooth unimodal peak structure becomes a multi-modal structure as shown in Figs.6(d) and 6(e). That interference effect nearly disappears at $t\approx 4\tau.$
@@ IV.2.1 Approximate expression for longitudinal momentum distribution function
In this subsection, we aim to analyze the oscillatory behavior of the LMS in the late-time limit, specifically for $t>2\tau$. We utilize the previously derived approximate analytical expression for the distribution function Eq.(98). By setting $p_{\perp}=0$, we derive an approximate expression that depends solely on the longitudinal momentum and the transformed time variable $y$. The longitudinal momentum distribution function $f(p_{\parallel},y)$ is then expressed as a series expansion in powers of $(1-y)$ up to the second order and neglecting higher-order terms:
@@
Following Eqs. (99), (100), and (101), it is evident that for $t\gg 2\tau$, the coefficients of $(1-y)$ and $(1-y)^{2}$ can be further approximated by retaining only the dominant contributions and disregarding the others, as follows:
@@
$\mathrm{C}_{1}(p_{\parallel},y)$ $\approx-4y|\Gamma_{1}\overline{\Gamma_{2}}|\frac{E_{0}\tau}{\omega_{1}}\cos{(\Upsilon)},$ (111)
$\mathrm{C}_{2}(p_{\parallel},y)$ $\approx\frac{4E_{0}^{2}\tau^{2}(\omega_{1}-P_{1})}{\omega_{1}^{2}}|\Gamma_{1}|^{2}+4|\Gamma_{1}\overline{\Gamma_{2}}|\cos{(\Upsilon)}\Biggl{[}\omega_{0}(1+3P_{1}(\omega_{1}+P_{1}))-\omega_{1}(3P_{1}(1+\omega_{1})+1)$
@@
Figure 8: Time evolution of the first and second terms present in the longitudinal momentum distribution function $f(p_{\parallel},t)$ (see Eq. (109)) as a function of $p_{\parallel}$. Red curve: $(1-y)\mathrm{C}_{1}$ and Blue curve: $(1-y)^{2}\mathrm{C}_{2}$. The field parameters are $E_{0}=0.1E_{c}$ and $\tau=10\,[m^{-1}]$.
In Figure 8, we illustrate the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ and the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$. These terms are plotted as functions of longitudinal momentum at different times, allowing for a comparison of their behavior over time. Fig. 8(a), we observe that during earlier times, the term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ is more pronounced compared to $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$. The profile of $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ shows a prominent peak around $p_{\parallel}\simeq-2[m]$, largely influenced by $|\Gamma_{1}|^{2}$ (refer to Eq. (112)), which peaks near $p_{\parallel}\approx-E_{0}\tau$ depending on parameters $E_{0}$ and $\tau$. The oscillations observed within $-1<p_{\parallel}<1$ stem from the $\cos{(\Upsilon)}$ in the expression for $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$, contributing to this oscillatory pattern. Over time, the influence of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ gradually increases compared to earlier times. This term exhibits oscillations within a Gaussian envelope, with the oscillations decaying away from $p_{\parallel}\approx 0$ (see Figs. 8(b) and (c)). To understand the oscillations and their dependence on $p_{\parallel}$-value and the time variable $y$, the argument $\Upsilon$ of the cosine function is approximated as follows:
$\cos{(\Upsilon)}$ $=\cos{\left(K_{0}+(p_{\parallel}+E_{0}\tau)(K_{1}+K_{2}(p_{\parallel}+E_{0}\tau))\right)},$ (113)
@@
$\mathrm{A}_{C1}$ $=\frac{4|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau}{\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}}}.$ (117)
@@
This leads us to find that the maxima occur at approximately $p_{\parallel}\approx-0.6$. The function exhibits a nearly Gaussian-like structure.
@@
At $t=4\tau$, the magnitude of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ dominates over the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ as seen in Fig.8(c), and the second-order term diminishes at $t=5\tau$ (see Fig.8(d)). It’s worth noting that both orders have a $(1-y)$ factor, which contributes to the distribution function. However, this contribution gradually decreases as time progresses, becoming less important in the late-time limit $y\simeq 1$. At $t\gg 2\tau$, the onset of oscillations in the Gaussian-like structure can be explained by the zeroth and first-order terms of the power series expansion of $f(p_{\parallel},y)$.
In figure 10, the comparison of $\mathrm{C}_{1}(p_{\parallel},y)$, $\mathrm{C}_{0}(p_{\parallel},y)$, and $\mathrm{C}_{0}(p_{\parallel},y)+(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ is shown to identify which order is responsible for the oscillatory features of LMS at $t\geq 3\tau$. At $t=3\tau$, $\mathrm{C}_{0}(p_{\parallel},y)$ exhibits a smooth single peak located at $p_{\parallel}=0$, while $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ shows oscillatory behavior within the small range $-1<p_{\parallel}<1$. The combined behavior resembles a smooth Gaussian structure with onset oscillations contributing to the distribution function, as confirmed by Fig. 10(a). As time progresses, the oscillation amplitude decreases due to the presence of the $(1-y)$ factor in the first-order term, reducing its impact (see Fig. 10(b)). Eventually, $\mathrm{C}_{0}(p_{\parallel},y)$ dominates and explains the distribution function’s behavior at asymptotic times as shown in Fig.10(c).
@@
On careful examination of the individual components of the longitudinal momentum distribution function in Eq.(109), we observe that during electron-positron formation at $t>2\tau$, the second-order term $\bigl{(}(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\bigr{)}$ is responsible for the primary peak located at $p_{\parallel}\approx-2$ (see Fig.5). As time progresses, this primary peak diminishes, and a secondary peak at $p_{\parallel}=0$ starts to build up due to the first-order term ($(1-y)\mathrm{C}_{1}(p_{\parallel},y)$) and zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y)$, which are responsible for the onset oscillation on that peak. As time progresses towards infinity, $\mathrm{C}_{1}(p_{\parallel},y)$ leads to suppression. Consequently, we observe only a secondary peak due to the dominance of the $\mathrm{C}_{0}(p_{\parallel},y)$ zeroth-order term( see Fig.5). Since the component $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ represents an oscillatory finite function whose magnitude depends on time, the magnitude of this function plays a crucial role in determining the dynamics of $f(p_{\parallel},t)$ in $p_{\parallel}$-space at finite times.
@@ IV.3 Vacuum Polarisation and depolarisation function
For a better understanding of the phenomenon of particle creation under a strong electric field, we will also trace the evolution of the vacuum polarization function $u(p_{\parallel},t),$ and its counterterm $v(p_{\parallel},t)$, which governs the depolarization. Figure 13 shows the time evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for different values of $p_{\parallel}$ at zero transverse momentum. Due to pair annihilation being stronger than pair creation in the early times, the depolarisation function $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$ (See Figs. 13(a) and (c)). It appears that in the polarization function, there is a sinusoidal function behavior. On the other hand, the depolarization function $v(p_{\parallel},t)$ shows an unimodal Gaussian peak structure in its temporal evolution. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ show oscillations with varying amplitudes during the initial transient stage. These oscillations are particularly pronounced when $p_{\parallel}=0[m]$, as evident from Figs.13(a) and (c). As time progresses, irregular oscillations are observed in the transient stage. However, the oscillations become regular and stable as the system enters the REPP stage. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ now exhibit regular oscillations centered around the zero value. Moreover, as the momentum value $p_{\parallel}$ increases, the amplitudes of these oscillations diminish, as shown in Fig. 13(b) and (d). During the REPP stage, one interesting finding is that $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ demonstrate balancing behavior characterized by similar oscillatory patterns. This balance results from the formation of real electron-positron pairs.
Figure 13: Evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for $p_{\parallel}=0$ (blue), $0.5$ (pink), $-0.5$ (brown), $1$ (red), and $-1$ (green) with $p_{\perp}=0$, $E_{0}=0.2E_{c}$, and $\tau=10\,[m^{-1}]$. All the measurements are taken in electron mass units.

### v6

@@ IV.2 Longitudinal Momentum Spectrum
From Figure 5, we see how particle creation proceeds. At early times, $t=-10[m^{-1}]$ in the QEPP region, when the electric field is increasing, the created particle shows a smooth unimodal-like structure of momentum spectrum with a peak at $p_{\parallel}\sim 2[m].$ The electric field propels the newly created particles towards the negative $z$-direction. The movement of the distribution function’s peak from $p_{\parallel}=2$ $(\penalty 10000\ eE_{0}\tau)$ to 2-2 $(\penalty 10000\ -eE_{0}\tau)$ can be understood through the concept of longitudinal quasi-momentum, $P(p_{\parallel},t)$ as depicted in the time evolution shown in Figs. 5(a) to (c). This momentum shift implies that the longitudinal momentum distribution of the generated particles is expected to span a range of $\Delta p_{\parallel}=2$, with the spread determined by the electric field parameters. At $t=0$ (electric field is maximum), the momentum distribution function, $f(p_{\parallel}=0,t=0)\approx 5\times 10^{-3}$, follows the tendency $E^{2}(t)/8$ Schmidt et al. (1999). For $t>0$, a reduction in the field strength with a decrease in an intriguing longitudinal momentum distribution function. It rapidly drops to $99.2\%$ of its maximum value, driven by its reliance on the strength of the electric field at $t=10[m^{-1}]$, as illustrated in Fig. 5(c). Within a narrow range of $1<p_{\parallel}<1$, the typically smooth unimodal profile of the longitudinal momentum spectra is disrupted, as depicted in Fig. 5(d).
@@
Certainly, we can identify specific time scales associated with the quantum signature that manifest as oscillations within the LMS of created pairs. Drawing from the emergence of a secondary peak, we define three distinct time scales: $(i)$ $t_{cp}$ (central peak formation): This time scale is characterized by the emergence of the secondary peak and the beginning of its development, $(ii)$ $t_{sep}$ (peak separation): At this time, the central peak becomes dominant or the time after the two peaks become distinctly separated, $(iii)$ $t_{dis}$ (disappearance of oscillation): This indicates the time when the oscillations within the central peak fade away or the time after the primary peak (or the left-side peak) ceases to exist. By identifying and quantifying these time scales, we can better characterize the intricate quantum dynamics reflected in the LMS during the REPP stage.
   $E_{0}\penalty 10000\ [E_{c}]$    $t_{cp}\penalty 10000\ [m^{-1}]$    $t_{sep}\penalty 10000\ [m^{-1}]$    $t_{dis.}\penalty 10000\ [m^{-1}]$
@@
Table 2: Effect of electric field strength $E_{0}$ on the time scales related to the formation of a central peak. $t_{cp}$ denotes the appearance of a central peak, $t_{sep}$ the time when two peaks become distinctly separated, and $t_{dis.}$ the disappearance of the oscillation in LMS.
@@
We have not yet discussed the Keldysh parameter, $\gamma$, which classifies the pair production process as either a multi-photon or tunneling mechanism. Additionally, we are investigating whether the oscillatory behavior explicitly depends on $\gamma$. For this study, we select two different combinations of $E_{0}$ and $\tau$ to vary $\gamma$ and examine its behavior over a finite time interval. Specifically, we set $\gamma=1$, corresponding to two distinct parameter configurations $(E_{0},\tau)$ associated with the Sauter pulse electric field. It is known that $\gamma=1$ represents an intermediate regime involving both tunneling and multi-photon processes, which are inherently non-perturbative Avetissian et al. (2002). This approach allows us to demonstrate how the same $\gamma$ value can lead to qualitatively different outcomes in the time evolution of the pair production process. Figure 6 shows LMS for the Sauter pulse electric field with $E_{0}=0.2E_{c}$ and $\tau=5\,[m^{-1}]$, while Fig.7 shows LMS for field parameter $E_{0}=0.1E_{c}$ and $\tau=10\,[m^{-1}]$. From Figs. 6 (a) to (c) we can see that due to the electric field smooth unimodal peaked profile accelerated towards negative $z$- direction and after the pulse duration $t=6[m^{-1}]$ we observed smooth profile becomes deformed near the region $0.5<p_{\parallel}<0.5.$ At $t=11.5[m^{-1}]$, the interference effect is observed, due to which the smooth unimodal peak structure becomes a multi-modal structure as shown in Figs.6(d) and 6(e). That interference effect nearly disappears at $t\approx 4\tau.$
@@ IV.2.1 Approximate expression for longitudinal momentum distribution function
In this subsection, we aim to analyze the oscillatory behavior of the LMS in the late-time limit, specifically for $t>2\tau$. We utilize the previously derived approximate analytical expression for the distribution function Eq.(98). By setting $p_{\perp}=0$, we derive an approximate expression that depends solely on the longitudinal momentum and the transformed time variable $y$. The longitudinal momentum distribution function $f(p_{\parallel},y)$ is then expressed as a series expansion in powers of (1y)(1-y) up to the second order and neglecting higher-order terms:
@@
Following Eqs. (99), (100), and (101), it is evident that for $t\gg 2\tau$, the coefficients of (1y)(1-y) and $(1-y)^{2}$ can be further approximated by retaining only the dominant contributions and disregarding the others, as follows:
@@
$\mathrm{C}_{1}(p_{\parallel},y)$ $\approx-4y|\Gamma_{1}\Gamma^{*}_{2}|\frac{E_{0}\tau}{\omega_{1}}\cos{(\Upsilon)},$ (111)
$\mathrm{C}_{2}(p_{\parallel},y)$ $\approx\frac{4E_{0}^{2}\tau^{2}(\omega_{1}-P_{1})}{\omega_{1}^{2}}|\Gamma_{1}|^{2}+4|\Gamma_{1}\Gamma^{*}_{2}|\cos{(\Upsilon)}\Biggl[\omega_{0}(1+3P_{1}(\omega_{1}+P_{1}))-\omega_{1}(3P_{1}(1+\omega_{1})+1)$
@@
Figure 8: Time evolution of the first and second terms present in the longitudinal momentum distribution function $f(p_{\parallel},t)$ (see Eq. (109)) as a function of $p_{\parallel}$. Red curve: $(1-y)\mathrm{C}_{1}$ and Blue curve: $(1-y)^{2}\mathrm{C}_{2}$. The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10\,[m^{-1}]$.
In Figure 8, we illustrate the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ and the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$. These terms are plotted as functions of longitudinal momentum at different times, allowing for a comparison of their behavior over time. Fig. 8(a), we observe that during earlier times, the term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ is more pronounced compared to $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$. The profile of $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ shows a prominent peak around $p_{\parallel}\simeq-2[m]$, largely influenced by $|\Gamma_{1}|^{2}$ (refer to Eq. (112)), which peaks near $p_{\parallel}\approx-E_{0}\tau$ depending on parameters $E_{0}$ and $\tau$. The oscillations observed within $1<p_{\parallel}<1$ stem from the $\cos{(\Upsilon)}$ in the expression for $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$, contributing to this oscillatory pattern. Over time, the influence of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ gradually increases compared to earlier times. This term exhibits oscillations within a Gaussian envelope, with the oscillations decaying away from $p_{\parallel}\approx 0$ (see Figs. 8(b) and (c)). To understand the oscillations and their dependence on $p_{\parallel}$ value and the time variable $y$, the argument $\Upsilon$ of the cosine function is approximated as follows:
$\cos{(\Upsilon)}$ $=\cos{\left[K_{0}+(p_{\parallel}+E_{0}\tau)(K_{1}+K_{2}(p_{\parallel}+E_{0}\tau))\right]},$ (113)
@@
$\mathrm{A}_{C1}$ $=\frac{4|\Gamma_{1}\Gamma^{*}_{2}|E_{0}\tau}{\sqrt{1+(p_{\parallel}+E_{0}\tau)^{2}}}.$ (117)
@@
This leads us to find that the maxima occur at $p_{\parallel}\approx-0.6$. The function exhibits a nearly Gaussian-like structure.
@@
At $t=4\tau$, the magnitude of the first-order term $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ dominates over the second-order term $(1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)$ as seen in Fig.8(c), and the second-order term diminishes at $t=5\tau$ (see Fig.8(d)). It’s worth noting that both orders have a (1y)(1-y) factor, which contributes to the distribution function. However, this contribution gradually decreases as time progresses, becoming less important in the late-time limit $y\simeq 1$. At $t\gg 2\tau$, the onset of oscillations in the Gaussian-like structure can be explained by the zeroth and first-order terms of the power series expansion of $f(p_{\parallel},y)$.
In figure 10, the comparison of $\mathrm{C}_{1}(p_{\parallel},y)$, $\mathrm{C}_{0}(p_{\parallel},y)$, and $\mathrm{C}_{0}(p_{\parallel},y)+(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ is shown to identify which order is responsible for the oscillatory features of LMS at $t\geq 3\tau$. At $t=3\tau$, $\mathrm{C}_{0}(p_{\parallel},y)$ exhibits a smooth single peak located at $p_{\parallel}=0$, while $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ shows oscillatory behavior within the small range $1<p_{\parallel}<1$. The combined behavior resembles a smooth Gaussian structure with onset oscillations contributing to the distribution function, as confirmed by Fig. 10(a). As time progresses, the oscillation amplitude decreases due to the presence of the (1y)(1-y) factor in the first-order term, reducing its impact (see Fig. 10(b)). Eventually, $\mathrm{C}_{0}(p_{\parallel},y)$ dominates and explains the distribution function’s behavior at asymptotic times as shown in Fig.10(c).
@@
On careful examination of the individual components of the longitudinal momentum distribution function in Eq.(109), we observe that during electron-positron formation at $t>2\tau$, the second-order term $\bigl((1-y)^{2}\mathrm{C}_{2}(p_{\parallel},y)\bigr)$ is responsible for the primary peak located at $p_{\parallel}\approx-2$ (see Fig.5). As time progresses, this primary peak diminishes, and a secondary peak at $p_{\parallel}=0$ starts to build up due to the first-order term ($(1-y)\mathrm{C}_{1}(p_{\parallel},y)$) and zeroth-order term $\mathrm{C}_{0}(p_{\parallel},y)$, which are responsible for the onset oscillation on that peak. As time progresses towards infinity, $\mathrm{C}_{1}(p_{\parallel},y)$ leads to suppression. Consequently, we observe only a secondary peak due to the dominance of the $\mathrm{C}_{0}(p_{\parallel},y)$ zeroth-order term( see Fig.5). Since the component $(1-y)\mathrm{C}_{1}(p_{\parallel},y)$ represents an oscillatory finite function whose magnitude depends on time, the magnitude of this function plays a crucial role in determining the dynamics of $f(p_{\parallel},t)$ in $p_{\parallel}$-space.
@@ IV.3 Vacuum Polarisation and depolarisation function
For a better understanding of the phenomenon of particle creation under a strong electric field, we will also trace the evolution of the vacuum polarization function $u(p_{\parallel},t)$ and its counterterm $v(p_{\parallel},t)$, which governs the depolarization. Figure 13 shows the time evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for different values of $p_{\parallel}$ at zero transverse momentum. Due to pair annihilation being stronger than pair creation in the early times, the depolarisation function $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$ (See Figs. 13(a) and (c)). It appears that in the polarization function, there is a sinusoidal function behavior. On the other hand, the depolarization function $v(p_{\parallel},t)$ shows an unimodal Gaussian peak structure in its temporal evolution. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ show oscillations with varying amplitudes during the initial transient stage. These oscillations are particularly pronounced when $p_{\parallel}=0[m]$, as evident from Figs.13(a) and (c). As time progresses, irregular oscillations are observed in the transient stage. However, the oscillations become regular and stable as the system enters the REPP stage. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ now exhibit regular oscillations centered around the zero value. Moreover, as the momentum value $p_{\parallel}$ increases, the amplitudes of these oscillations diminish, as shown in Fig. 13(b) and (d). During the REPP stage, one interesting finding is that $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ demonstrate balancing behavior characterized by similar oscillatory patterns. This balance results from the formation of real electron-positron pairs.
Figure 13: Evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for $p_{\parallel}=0$ (blue), $0.5$ (pink), 0.5-0.5 (brown), 11 (red), and 1-1 (green) with $p_{\perp}=0$, $E_{0}=0.2E_{c}$, and $\tau=10\,[m^{-1}]$. All the measurements are taken in electron mass units.